Shift report 20030215 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Feb 15 2003 - 00:02:36 EST

  • Next message: Apache: "Shift report 20030215 00:00-08:00"
    ------------------------------------------------------------
    Date: Sat Feb 15 00:02:35 2003, Shift: 16:00-24:00
    Supervisor: ZY, with: 
    Log book#14 pages 17 - 17
    
    Shift Summary:
    ------------------------------------------------------------
    1. Good store from 20:15 with initial ZDC rate of 5.5K.
    2. Data runs
       7897 with trig. 6 74k, trig. 3 279k  
       7899 with trig. 6 37k , trig. 3 105k
    3. Since T3-A1-yv tripped, according Pawel's instruction, voltage on A1-yv and C1-yv are set to 100V and -100V respectively.
       During run 7897, D5 went to standby at some point due to cooling fault. People from MCR said it might be around 21:00.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Feb 15 2003 - 00:06:30 EST